OpenStudy (anonymous):

could anyone help me?!? :D find the inverse function of g(x)=x^2+2x+3 when x is greater than or equal to -1.

jimthompson5910 (jim_thompson5910):

are you able to complete the square for that quadratic?

OpenStudy (anonymous):

is there a way to solve it without completing the square? cuz i dont really remember how to do that haha...been a long summer...

jimthompson5910 (jim_thompson5910):

no not that I can see

OpenStudy (anonymous):

oh could u refresh me on how to do it then ? :/

jimthompson5910 (jim_thompson5910):

focus on x^2+2x

jimthompson5910 (jim_thompson5910):

what term is missing to make that a perfect square?

OpenStudy (anonymous):

+1?

OpenStudy (anonymous):

x^2+2x+1?

OpenStudy (anonymous):

:(

jimthompson5910 (jim_thompson5910):

yes that is correct

jimthompson5910 (jim_thompson5910):

so if we had x^2 + 2x + 1, we could factor that to (x+1)^2

jimthompson5910 (jim_thompson5910):

however, we have x^2+2x+3

jimthompson5910 (jim_thompson5910):

one thing we can do is break up that 3 into 1+2 so we now have x^2+2x+1+2

jimthompson5910 (jim_thompson5910):

then we factor x^2 + 2x + 1 to get (x+1)^2 so we now have g(x) = (x+1)^2 + 2

jimthompson5910 (jim_thompson5910):

making sense?

OpenStudy (anonymous):

yes so far

jimthompson5910 (jim_thompson5910):

so we have y = (x+1)^2 + 2

jimthompson5910 (jim_thompson5910):

how do we find the inverse of that?

OpenStudy (anonymous):

change y and x and solve for y again?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so solve for y in x = (y+1)^2 + 2

OpenStudy (anonymous):

\[y=+/- \sqrt{x-2}-1\]

OpenStudy (anonymous):

ok, so why was the domain restricted to begin with?

jimthompson5910 (jim_thompson5910):

or \[\Large y=-1 \pm \sqrt{x-2}\] but you pretty much got it. There's one slight issue though

jimthompson5910 (jim_thompson5910):

let's look at a rough graph of y = x^2+2x+3

jimthompson5910 (jim_thompson5910):

That is the same as y = (x+1)^2 + 2 so we see that the vertex is (-1,2) since that equation above is in vertex form y = a(x-h)^2 + k

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

Rough Sketch of y = x^2+2x+3 or y = (x+1)^2 + 2